Order: Vancomycin 1g ivpb q12h in 100 ml of normal saline (0.9% NaCl). At how many ml/hr will the nurse set the pump to deliver the vancomycin over 30 minutes? _______________mL/hr

The clinician orders 1.5 liter of fluids to infuse in 16 hours. You have available administration set with a drop factor of 10 gtt/mL. How many gtt/min should the fluids infuse? ___________ gtt/min

  • The provider ordered Diazepam 7.5 mg IM stat for a client with agitation. How many mL will you administer to the client? Diazepam comes in 5mg/mL preparation. _________mL
  • Order: Heparin 3,500 units SQ every 6 hours. Available: Vial containing 5,000 units/mL. How many mL will the client receive? ______________ mL
  • Order: Lovenox 30mg SQ Q12H after hip replacement. The pharmacy has sent 40mg/0.4mL in a prefilled syringe. How many mL will be administered? ______________ mL
  • A physician orders “Cloxacillin 0.5g PO q12 hours”. Cloxacillin comes in 500mg preparation. How many tablet/s will be needed at each dose to administer the correct dose? ________tablet/s
Answer #1

ANSWER.

* Calculation of ml/hr.

         ml/hr = Volume(ml) / Time(hr).

Here

      Volume(ml) = 100 ml.

       Time(hr) = 0.5 hr (30 min).

As per above formula

           ml/hr = 100 / 0.5

                    = 200 ml/hr.

* Calculation of gtt/min.

         Formula

              gtt/min = Volume(ml) / Time(min) * Drop factor(gtt/ml)

    Here

         Volume(ml) = 1500 ml (1.5l)

          Time(min) = 960 min (16 hrs)

          Drop factor = 10 gtt/ml.

As per above formula

           gtt/min = 1500 / 960 * 10

                      = 1.56 * 10

                      = 15.6 gtt/min.

* Calculation of volume(ml) of drug required.

         Formula

             Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

          Desired dose(D) = 7.5 mg.

          Dose in hand (H) = 5 mg.

               Quantity(Q) = 1 ml

    As per above formula

        Volume(ml) of drug required = D/H *Q

                                                      = 7.5/5 * 1

                                                      = 1.5 * 1

                                                      = 1.5 ml.

* Calculation of volume(ml) of drug required .

        Formula

            Volume(ml) = Desired dose / Dose in hand * Quantity.

    Here

        Desired dose (D) = 3500 unit.

        Dose in hand (H) = 5000 unit.

            Quantity (Q) = 1 ml

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 3500/5000 * 1

                                                     = 0.7 * 1

                                                     = 0.7 ml.

* Calculation of volume(ml) of drug required.

        Formula

            Volume(ml) = Desired dose/Dose in hand * Quantity.

     Here

         Desired dose (D) = 30 mg.

         Dose in hand (H) = 40 mg.

              Quantity(Q) = 0.4 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 30/40 * 0.4

                                                      = 0.75 * 0.4

                                                      = 0.3 ml.

* Calculation of number of tablets required for each dose.

           Ordered dose = 500 mg (0.5 g) Q12H

           Available dose = 500 mg/tablet.

    So

        Number of tablets required per each dose = Ordered dose / Available dose.

                                                                            = 500 / 500

                                                                            = 1 tablet per dose.